The mean lifetime of a sample of 144 fluorescent light bulbs produced by a company is computed to be 170 hours with standard deviation of 120 hours.Test the company's claim that its bulbs' mean time is 200 hours at 0.05 significance level
1
Expert's answer
2021-09-17T03:49:51-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=200"

"H_1:\mu\not=200"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=144-1=143" degrees of freedom, and the critical value for a two-tailed test is "t_c=1.976692."

The rejection region for this two-tailed test is "R=\{t:|t|>1.976692\}."

The t-statistic is computed as follows:


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{170-200}{120/\sqrt{144}}=-3"

Since it is observed that "|t| = 3 > t_c = 1.976692=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for two-tailed "\alpha=0.05, df=143,t=-3" is "p=0.003186," and since "p=0.003186<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 200, at the "\alpha=0.05" significance level.